A mississippi riverboat can travel 30km downstream in 3 hours and make the return trip in 5 hours. Find the speed of the boat in still water.
Answer by stanbon(75887)

A mississippi riverboat can travel 30km downstream in 3 hours and make the return trip in 5 hours. Find the speed of the boat in still water.
------------
Downstream DATA:
distance = 30 km ; time = 3 hrs ; rate = d/t = 10kph
---------------
Upstream DATA:
distance = 30 km ; time = 5 hrs ; rate = d/t = 6 kph
----------------
let b be the speed of the riverboat in still water
let c be the speed of the current
------------------------
EQUATION:
b+c=10 kph
b-c=6 kph
-------------
Add to get:
2b = 16 kph
b = 8 kph (this is the speed of the riverboat in still water)
---------------
Substitute to solve for current speed:
8+c=10
c=2 kph (this is the speed of the current)
================
Cheers,
Stan H.